OpenStudy (anonymous):

Determine which value is a solution to the open sentence. 3 + 6x = 21 A. 18 B. 4 C. 3 D. 12

OpenStudy (anonymous):

subtract 3 from 21 then divide that answer by 6

OpenStudy (anonymous):

^ she is absoltely correct:) The answer is simple

OpenStudy (anonymous):

i dont understand cuz there is no 6 in the answer

OpenStudy (anonymous):

the answer isn't 6 ill show you the steps

OpenStudy (anonymous):

yes i know. But there is one in the problem 21-3=18 18 divided by 6 = 4 There fore the answer is 4

OpenStudy (anonymous):

3+6x=21 -3 -3 6x=18 now just take 18 divided by 6 and youll have ur answer

OpenStudy (anonymous):

i mena 3 my bad

OpenStudy (anonymous):

mean

OpenStudy (anonymous):

18/6 is actually 3 @Littlemsadventuretime

OpenStudy (anonymous):

yeah i got that lol

OpenStudy (anonymous):

ya I seen that I posted it right after

OpenStudy (anonymous):

thanks you :)

OpenStudy (anonymous):

No problem. Need anymore help?

OpenStudy (anonymous):

let me see :)

OpenStudy (anonymous):

If so post it as another question. And tag my name:)

OpenStudy (anonymous):

how do i tag your name

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

i need help but its not math :/

OpenStudy (anonymous):

@crotin1111

OpenStudy (anonymous):

Ok post it in the other subject, ill try to help:)